A manager uses the following equation below to forecast quarterly demand for a product: Yt =6,000+...

  1. A manager uses the following equation below to forecast quarterly demand for a product:

Yt =6,000+ 50 t, where t = 0 at Q2 of last year (2010). Quarter relatives are Q1 = .6, Q2 = .9, Q3= 1.1, and Q4=1.2. What forecasts are appropriate for the last quarter (Q4) of this year (2011)?

  1. 5,400
  2. 6,060
  3. 7,440
  4. 7,560
  5. None of the above
Answer #1

The trend equation for monthly demand: Ft = 6000 + 50t

Where t = 0 in Q2 of last year. In this sequence the value of t in Q4 of this year is calculated below

t= 0 (Q2 last year)

t= 1 (Q3 of last year)

t = 2 ( Q4 of last year)

t = 3 (Q1 of this year)

t= 4 (Q2 of this year)

t = 5 (Q3 of this year)

t = 6 (Q4 of this year)

So using the trend equation unadjusted forecast for Q4 of this year = 6000+50t = 6000+(50×6) = 6000+300 = 6300

Trend adjusted Forecast for Q4 of this year = unadjusted forecast × Seasonal Relative = 6300 × 1.2 = 7560

so the answer is option D